OpenStudy (anonymous):

Solve the ff. Linear Equation and check 1.)4x-2=6x-12 2.)5+2(3-x)=4+2(x-2)+5x 3.)(x-2)/3=3-(x-9)/3 4.)x^2+7x+10=x^2+5x-6 5.)1/2(3x+5)=2x-3

Parth (parthkohli):

Hey, welcome to OpenStudy! Can you let me know where you have the problem?

OpenStudy (anonymous):

solve the ff Linear equations

Parth (parthkohli):

No, I mean can you tell me which question you are specifically having trouble with?

OpenStudy (anonymous):

all

Parth (parthkohli):

I'd help you with the first to start with. Add two to both sides. \( \color{Black}{\Rightarrow 4x - 2 + 2 = 6x - 12 + 2 }\) \( \color{Black}{\Rightarrow 4x +0 = 6x - 10 }\) \( \color{Black}{\Rightarrow 4x = 6x -10}\) Next step is to subtract 6x from both sides. \( \color{Black}{\Rightarrow 4x - 6x = 6x - 6x - 10 }\) \( \color{Black}{\Rightarrow -2x = 0 - 10}\) \( \color{Black}{\Rightarrow -2x = -10}\) Now, divide both sides by -2. \( \color{Black}{\Rightarrow x = -10 \div -2 }\)

Parth (parthkohli):

-10/-2 = 5 right?

OpenStudy (anonymous):

yeah SMART :))

Parth (parthkohli):

Can you do the 2nd for me?

OpenStudy (anonymous):

number 1 is only the easy one :)) can you please do all ?

Parth (parthkohli):

Can you please post it in a new one?

OpenStudy (anonymous):

okay